OpenStudy (anonymous):

Can someone help explain how to do this question. I am having great difficulty working it out. :P If the steel cables on an aircraft carrier stops a plane in 150m with an acceleration of -15m/s^2, find the time the plane takes to stop

OpenStudy (loser66):

you have formula \[ V^2 = V_0^2 + 2as\] at stop, V =0 a = -15m/s^2 therefore \(V_0 = 2as = \sqrt{15*150} = 47.43m/s\) apply \[x = x_0 +v_0t +\frac{1}{2}at^2\] where \(x_0 =0\) you have time t as \[\frac{15}{2}t^2-47.43t+150=0\] solve for t =6.33s

OpenStudy (anonymous):

Oh thank you so much.